Week of 180122

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Andrea (MW Officer + data management), Borja (monitoring), Gavin (computing), Julia (WLCG), Maarten (SCOD + ALICE), Marian (networks + monitoring), Vincent (security)
  • remote: Darren (RAL), David B (IN2P3), David M (FNAL), Di (TRIUMF), Dmytro (NDGF), Kyle (OSG), Marcelo (CNAF), Onno (NLT1), Peter (ATLAS), Renato (LHCb), Stephan (CMS), Victor (JINR), Vincenzo (EGI), Xavier (KIT), Xin (BNL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Frontier bug encountered (old bug from 2015) where squids serve a stale error result, workaround is to flush squid cache. Bug being handled FTAPPDEVEL-132
    • Case of T2 site running DPM on SL5, offline/unresponsive pending upgrade, being handled now but should remind all T2 about MoU committments.

  • CMS reports ( raw view) -
    • Mo major issues
    • Uneventful recovery from the Monday Spectre/Meltdown reboots, not looking forward to this weeks reboots
    • Smooth running at about 200k cores (1/4 analysis, 3/4 production)
    • Some xrootd issues where redirectors needed a kick

  • ALICE -
    • Normal activity level on average
    • Grid got drained on Wed due to central service issue
      and on Fri because of related maintenance intervention.

  • LHCb reports ( raw view) -
    • Activity
      • HLT farm "returning" from cooling maintenance.(no jobs running yet)
      • 2016 data restripping running full steam. Almost all data processed (waiting for CNAF)
      • Monte Carlo productions using remaining resources.
    • Meltdown & Spectre, several voboxes rebooting this week.
    • Site Issues
      • CERN/T0
        • NTR
      • T1
        • GRIDKA problems with FTS transfers(from and to) and "put and register".
          • Xavier: the issue has been resolved;
            all dCache pools available to LHCb were full,
            but the space reservation was not shown as full;
            extra space has been added to overcome that

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • CNAF: Starting to power up and setting up critical INFN services.
  • EGI: NTR
  • FNAL: NTR
  • IN2P3: NTR
  • JINR: NTR
  • KISTI:
  • KIT:
    • Xavier: besides the issue reported by LHCb we also are looking into
      importing SAM results into our local monitoring and ran into an
      unexpected matter for LHCb (SE tests not in critical profile) and
      a difficulty for ALICE (some hostnames absent in JSON)
    • Maarten: please open GGUS tickets for VO Support about these issues
  • NDGF: NTR
  • NL-T1: NTR
  • NRC-KI:
  • OSG: NTR
  • PIC:
  • RAL: NTR
  • TRIUMF: NTR

  • CERN site:
    • The remaining service availability zones will be rebooted Mon, Tue, Wed this week OTG0041682. Please monitor OTG, as the schedule may change.
  • CERN computing services:
    • CERN batch service is currently half-draining, for the first reboots on Thursday 25th. The 2nd half will then be drained for reboot on Tuesday 30th. While draining, jobs whose wall-clock fits will be accepted.
  • CERN storage services:
    • tomorrow EOS-CMS will be upgraded to the Citrine release
    • all FTS nodes have been rebooted OK in the meantime
  • CERN databases:
  • GGUS: NTR
  • Monitoring:
    • Final reports for the December availability sent around
  • MW Officer: NTR
  • Networks: NTR
  • Security: Spectre/Meltdown update:
    • RedHat & derivative removed microcode updates ("instabilities")
    • Recommendations:
      • Update kernel & reboot as soon as possible (to fix Meltdown & Spectre variant 1)
      • Stay tuned for updates on microcode for Spectre variant 2 (pending Intel releases...)

AOB:

Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2018-01-22 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback